OpenStudy (anonymous):

Write the equation so that y is the function of x. 14-6x=7y. please show all steps.

OpenStudy (anonymous):

Divide by 7 to get y=14-6x/7

OpenStudy (turingtest):

divide both sides by 7 2-(6/7)x=y

OpenStudy (anonymous):

I have to rewrite the equation too.

OpenStudy (turingtest):

ndog forgot to divide the 14 so his is wrong

OpenStudy (anonymous):

Can you rewrite the equation too so that y is the function of x.

OpenStudy (anonymous):

I guess it's y=-6x/7+2

OpenStudy (turingtest):

yup

OpenStudy (turingtest):

but you should see that y=-6x/7+2 is the same as 2-(6/7)x=y they are both y as a function of x, whether you put the y on the left or right. All that matters is that y is all alone on one side of the = sign.